# Blowing star formation away in AGN Hosts (BAH) - II. Investigating the origin of the H2 emission excess in nearby galaxies with JWST MIRI

### v2

@@ 1 Introduction
The use of JWST, in combination with the Mid-Infrared Instrument (MIRI) medium-resolution spectrometer (MRS), enables detailed studies of the emission and dynamics of warm molecular gas in nearby galaxies, utilizing the MIR rotational H2 lines. Results obtained with MIRI in the study of molecular gas in nearby galaxies include: increased emissions of warm and hot H2 as well as ionized gas, which coincide with the intersection of the jet with dust lanes in the disk of NGC 7319 (Pereira-Santaella et al., 2022); regions of increased H2 velocity dispersion between the nucleus and the star-forming ring in NGC 7469, attributed to shocks generated by ionized outflows (U et al., 2022); a wealth of interacting structures among various gas phases in the intergalactic medium in Stephan’s Quintet, with strong MIR H2 line emissions from warm gas generated by the fragmentation of dense cold molecular clouds, followed by mixing in the post-shock gas (Appleton et al., 2023); a collimated outflow of warm molecular gas originating from the southern nucleus of the ultra-luminous infrared galaxy NGC 3256, which is co-spatial with the outflow in cold gas observed by ALMA (Sakamoto et al., 2014). The warm to cold mass fraction is roughly 4%; a large fraction of warm molecular gas, accounting for 75% of the total molecular gas mass, in M83 (Hernandez et al., 2023); powerful jet-driven outflows in warm molecular gas in the inner few kpc of nearby () radio-loud AGN (Leftley et al., 2024; Costa-Souza et al., 2024); evidence of shock excitation of H2 and detection of warm molecular outflows within the ionization cones of the nearby Seyfert galaxy NGC 5728 (Davies et al., 2024).
@@ 2.1 The galaxies
In the optical, 3C 293 presents dust filaments spanning scales from 102–103 pc, which have been interpreted as a result of a previous merger event (e.g. Martel et al., 1999). Outflows co-spatial with the inner radio jet have been observed in ionized gas (Emonts et al., 2005; Mahony et al., 2016; Riffel et al., 2023a) with velocities of 1000 km s-1, and in neutral hydrogen (Morganti et al., 2003) with velocities of up to 1400 km s-1. The ionized gas emission in the inner 2 kpc of this galaxy is primarily attributed to photoionization by the central AGN, with an additional contribution from shocks for the outflow component, due to the interaction of the radio jet with the surrounding gas (Riffel et al., 2023a). While the outflows are not detected in cold molecular gas (Labiano et al., 2014), Costa-Souza et al. (2024) report the detection of outflows in warm H2, using JWST MIRI MRS data. The warm molecular outflows are observed co-spatial with the radio core and these authors estimated a mass-outflow rate of 6 M⊙ yr-1, which is one order of magnitude larger than the ionized gas outflows rate. If sustained, these outflows are capable of expelling all warm molecular gas from the central region of the galaxy in just 1 Myr (Costa-Souza et al., 2024).
@@ 4 Discussion
The plots of vs. log H2 S(3)/PAH11.3μm shown in Figs. 2, 3 and 4 clearly present two trends for all galaxies: (i) one group of points where the H2 S(3)/PAH11.3μm ratio increases with , and (ii) another group of points where the ratio remains practically constant. A correlation between the with H2 S(3)/PAH11.3μm may indicate a kinematic evidence that shocks are producing the H2 emission excess observed in these galaxies. To roughly distinguish between the two groups of points and spatially locate them within the galaxies, we follow Riffel et al. (2023b) and identify the kinematically disturbed regions (KDRs) and the virially dominated region (VDRs), using the measurements. When gas motion is primarily driven by the gravitational potential, the velocity dispersion is expected to decline with increasing distance from the nucleus. Consequently, the velocity dispersion observed at the nucleus can be interpreted as the upper limit of what is attributable to the gravitational potential. We establish a threshold to define the KDR, based on the fitting of the H2 S(3) emission-line nuclear profile. For each galaxy, we model the H2 S(3) emission line in the nuclear spaxel using two Gaussian components: a narrow component attributed to gas in the plane of the disc, and a broad component associated with outflows. The KDR correspond to locations with values larger than that of the nuclear narrow component, while other locations corresponds to the VDR, defined by spaxels with values smaller than the nuclear one. The limits are 370 km s-1 for 3C 293, 290 km s-1 for NGC 3884 and 385 km s-1 for CGCG 012-070.

### v4

@@ 1 Introduction
The use of JWST, in combination with the Mid-Infrared Instrument (MIRI) medium-resolution spectrometer (MRS), enables detailed studies of the emission and dynamics of warm molecular gas in nearby galaxies, utilizing the MIR rotational H2 lines. Results obtained with MIRI in the study of molecular gas in nearby galaxies include: increased emissions of warm and hot H2 as well as ionized gas, which coincide with the intersection of the jet with dust lanes in the disk of NGC 7319 (Pereira-Santaella et al., 2022); regions of increased H2 velocity dispersion between the nucleus and the star-forming ring in NGC 7469, attributed to shocks generated by ionized outflows (U et al., 2022); a wealth of interacting structures among various gas phases in the intergalactic medium in Stephan’s Quintet, with strong MIR H2 line emissions from warm gas generated by the fragmentation of dense cold molecular clouds, followed by mixing in the post-shock gas (Appleton et al., 2023); a collimated outflow of warm molecular gas originating from the southern nucleus of the ultra-luminous infrared galaxy NGC 3256, which is co-spatial with the outflow in cold gas observed by ALMA (Sakamoto et al., 2014). The warm to cold mass fraction is roughly 4%; a large fraction of warm molecular gas, accounting for 75% of the total molecular gas mass, in M83 (Hernandez et al., 2023); powerful jet-driven outflows in warm molecular gas in the inner few kpc of nearby () radio-loud AGN (Leftley et al., 2024; Costa-Souza et al., 2024); evidence of shock excitation of H2 and detection of warm molecular outflows within the ionization cones of the nearby Seyfert galaxy NGC 5728 (Davies et al., 2024). Ogle et al. (2025) used MIRI, along with near-infrared and optical integral field spectroscopy, to study Cygnus A, revealing a density-stratified NLR shaped by the interaction of the radio jet with the interstellar medium, and multi-phase biconical outflows reaching velocities of 600–2000 km s-1.
@@ 2.1 The galaxies
In the optical, 3C 293 presents dust filaments spanning scales from 102–103 pc, which have been interpreted as a result of a previous merger event (e.g. Martel et al., 1999). Outflows co-spatial with the inner radio jet have been observed in ionized gas (Emonts et al., 2005; Mahony et al., 2016; Riffel et al., 2023a) with velocities of 1000 km s-1, and in neutral hydrogen (Morganti et al., 2003) with velocities of up to 1400 km s-1. The ionized gas emission in the inner 2 kpc of this galaxy is primarily attributed to photoionization by the central AGN, with an additional contribution from shocks for the outflow component, due to the interaction of the radio jet with the surrounding gas (Riffel et al., 2023a). Previous Spitzer observations of 3C 293 reveal strong jet-shocked H2 emission (Ogle et al., 2010), indicating that the radio jet drives both the outflow and the heating of the warm H2 gas (Guillard et al., 2012). While the outflows are not detected in cold molecular gas (Labiano et al., 2014), Costa-Souza et al. (2024) report the detection of outflows in warm H2, using JWST MIRI MRS data. The warm molecular outflows are observed co-spatial with the radio core and these authors estimated a mass-outflow rate of 6 M⊙ yr-1, which is one order of magnitude larger than the ionized gas outflows rate. If sustained, these outflows are capable of expelling all warm molecular gas from the central region of the galaxy in just 1 Myr (Costa-Souza et al., 2024).
@@ 4 Discussion
The plots of vs. log H2 S(3)/PAH11.3μm shown in Figs. 2, 3 and 4 clearly present two trends for all galaxies: (i) one group of points where the H2 S(3)/PAH11.3μm ratio increases with , and (ii) another group of points where the ratio remains practically constant. A correlation between the with H2 S(3)/PAH11.3μm may indicate a kinematic evidence that shocks are producing the H2 emission excess observed in these galaxies. To roughly distinguish between the two groups of points and spatially locate them within the galaxies, we follow Riffel et al. (2023b) and identify the kinematically disturbed regions (KDRs) and the virially dominated region (VDRs), using the measurements. When gas motion is primarily driven by the gravitational potential, the velocity dispersion is expected to decline with increasing distance from the nucleus. Consequently, the velocity dispersion observed at the nucleus can be interpreted as the upper limit of what is attributable to the gravitational potential. We establish a threshold to define the KDR, based on the fitting of the H2 S(3) emission-line nuclear profile. For each galaxy, we model the H2 S(3) emission line in the nuclear spaxel using two Gaussian components: a narrow component attributed to gas in the plane of the disc, and a broad component associated with outflows. The KDR correspond to locations with values larger than that of the nuclear narrow component, while other locations corresponds to the VDR, defined by spaxels with values smaller than the nuclear one. The limits are 370 km s-1 for 3C 293, 290 km s-1 for NGC 3884 and 385 km s-1 for CGCG 012-070. The uncertainties in the values of the nuclear narrow component are less than 20 km s-1 for all three galaxies, resulting in an effect of less than 5 % on the fraction of spaxels that define the KDR in all objects. As discussed in Riffel et al. (2023b), for nearby AGN hosts, the different assumptions used to select spaxels and calculate the mass of the outflow gas lead to discrepancies of up to 1 dex in the ionized gas outflow rates between different methods, with approximately 0.3 dex of this discrepancy due to the limits used to define the spaxels dominated by outflows.
The cuts are below the escape velocity for all three galaxies. Using SDSS-based stellar masses (Maraston et al., 2009) and isophotal radii (Adelman-McCarthy et al., 2008), and assuming that the stellar mass corresponds to 15% of the total mass of the galaxies, we obtain rough escape velocities of 460 km s-1 for 3C 293, 540 km s-1 for NGC 3884, and 445 km s-1 for CGCG 012-070, estimated using a simple Keplerian approximation. For all three galaxies, the limits used to separate the KDR from the VDR are lower than the escape velocities. In 3C 293, 26% of the KDR spaxels have velocities above the escape velocity, which accounts for 51% of the H2 S(3) emission line flux in the KDR. None of the spaxels in NGC 3884 exceed the escape velocity, while 62% of the KDR spaxels in CGCG 012-070 have velocities above the escape velocity, accounting for 33% of the H2 S(3) KDR emission.